Can you help me understand how to approach this sequence:
3, -6, 12, 4, 20, ?

Answer by Edwin McCravy(20054)

3, -6, 12, 4, 20,

Look at the red numbers below.  They go -9,-8,-7,-6,...
Look at the blue numbers below. They go 18,16,14,12,... 

3 + -9 = -6
         -6 + 18 = 12
                   12 + -8 = 4
                             4 + 16 = 20
                                      20 + -7 = 13
                                                13 + 14 = 27
                                                          27 + -6 = 21
                                                                    21 + 12  

Notice also that each blue number is -2 times the preceding red number.

Edwin